OpenStudy (anonymous):

Two forces with magnitudes of 150 and 75 pounds act on an object at angles of 30° and 150° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

i'm really confused could someone help me i haven't done this before so i don't know how it works.

OpenStudy (anonymous):

Use the cosine rule. do you know it?

OpenStudy (anonymous):

yes but im not sure how to apply it

OpenStudy (anonymous):

first how do you get the direction?

OpenStudy (nurali):

F = 150 (cos(30), sin(30)) + 75 (cos(150), sin(150)) F = 150 (√3/2, 1/2) + 100 (-√3/2, 1/2) F = (75√3−50√3, 75+50) Magnitude of resultant force = √((75√3−50√3)² + (75+50)²) = 50√7 ≈ 132.287 pounds Direction of resultant force = tan⁻¹((75+50)/(75√3−50√3)) ≈ 70.89°

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thx! that helps so much:) i was on track with about the first step but i didn't know what to do after.

OpenStudy (nurali):

My Pleasure.